OpenStudy (anonymous):

6 friends (Andy, Bandy, Candy, Dandy, Endy and Fandy) are out to dinner. They will be seated in a circular table (with 6 seats). How many ways are there to seat them?

OpenStudy (anonymous):

\[6 \times 6 = 36\]

OpenStudy (anonymous):

yes the answer is 36. @wmckinely

OpenStudy (raden):

P = (n - 1)! put n = 6 to get P = (6 - 1)! = 5! = 120 ways

OpenStudy (anonymous):

yes The ans is 120

OpenStudy (anonymous):

120 is the answer according to each person's view.